### v3

# Structural dynamics and optimal transport of an active polymer

Abstract
We study the spontaneous configuration transitions of an active semi-flexible polymer between spiral and non-spiral states, and show that the configuration dynamics is fully described by a subcritical pitchfork bifurcation. Exploiting the fact that active polymer barely moves in spiral states and exhibits net displacements in non-spiral states, we theoretically prove that the motion of the active polymer is consistent with a run-and-tumble-like dynamics. Moreover, we find that there exists an optimal self-propelling force, at which the probabilities of finding the polymer in the spiral and non-spiral state become equal, that maximizes the diffusion coefficient.
Active polymers have recently become an interesting field of research in soft matter and non-equilibrium statistical physics. They play an important role in biological systems such as duplication of DNA by DNA-polymerase albert , synthesizes of proteins by ribosomes albert ; zia , chromatin in eukaryotic cells chromatin1 ; chromatin2 ; chromatin3 ; chromatin4 ; chromatin5 ; chromatin6 , movement of actin filaments by motor proteins in cellular cytoskeletons cytos1 ; cytos2 ; turning6 ; turning7 ; Motility_Assay_Book , microtubule bundles and spools in a kinesin-1 gliding motility assay Chaudhuri_Nano_Lett_2018 , sensing mechanical deformation of the soft materials surface by tracking the gluiding assay of microtubules as active probes on a carpet of kinesin-1 motor proteins Inoue_Nat_Commu_2015 , and the oscillatory beating to drive of flagella flagella0 ; flagella1 ; flagella2 . In all of the above examples, as mechanical tangential forces are generated on the backbone of the objects, they can be considered as active polymers. Motivated by these applications, many studies have focused, both experimentally exp1 ; exp2 ; exp3 ; exp4 ; exp5 and theoretically theo1 ; theo2 ; theo3 ; theo4 ; theo5 ; theo6 ; theo7 ; theo8 ; theo9 ; PRL ; theo10 ; baskaran ; theo11 ; turning ; theo12 ; Jabbari_PRE ; Jabbari_JCP , on active polymers. Theoretical studies have shown that the diffusion coefficient of the flexible/semi-flexible active polymer is a monotonically increasing function – linear or quadratic – of Péclet number theo1 ; theo5 ; theo6 ; theo7 ; theo9 ; PRL ; baskaran ; theo11 ; theo12 .
The structure and configuration of an active polymer have been studied in theo4 ; theo5 ; theo6 ; theo9 ; PRL ; theo10 ; baskaran ; turning ; theo11 ; Jabbari_PRE ; Jabbari_JCP . The interplay between thermal fluctuations and self-propelling stresses (i.e. activity) theo9 ; baskaran ; turning – as occurs when a polymer is embedded in a gliding assay with motor proteins as the source of activity Duke_PRL ; Bourdieu_PRE ; Bourdieu_PRL ; Grill_Nano_Lett ; Tas_Nano_Lett ; Farkas_book ; theo10 ; turning – leads to transient polymer configurations, classified into two main groups: spiral and non-spiral ones. For the latter case turning the radius of gyration (RG) is a convex function with respect to the attachment/detachment rates and Péclet number of the motor proteins, while for a system composed of an active polymer wherein the force is applied to all beads of the polymer theo9 RG is a monotonic decreasing function of the polymer activity [see the Supplemental Material (SM) for more details].
Figure 1: (a) Typical snapshot of a semi-flexible polymer with contour length of $N=200$ and persistence length of $\ell_{\textrm{p}}=5$ after equilibration and just before acting the self-propelling force $F_{\textrm{sp}}$ (orange arrow) on each segment. The polymer is divided to $N/\ell_{\textrm{p}}$ segments. The red and green beads represent the tail and head of each segment, respectively. In each segment $F_{\textrm{sp}}$ acts on the head monomer directing from tail to head (orange dashed line). (b) Definition of the spiral and non-spiral states: probability distribution function of the turning number $P(\psi)$ for fixed values of $F_{\textrm{sp}}=2$, $N=200$ and $\ell_{\textrm{p}}=5$. The vertical black dashed lines at minima of the $P(\psi)$ (located at the left $\psi_{\textrm{min}}^{\textrm{l}}$ and the right $\psi_{\textrm{min}}^{\textrm{r}}$ minima) separate the two spiral and non-spiral states. The polymer with $\psi_{\textrm{min}}^{\textrm{l}}<\psi<\psi_{\textrm{min}}^{\textrm{r}}$ is in the non-spiral state, and it is in the spiral state for $\psi<\psi_{\textrm{min}}^{\textrm{l}}$ or $\psi>\psi_{\textrm{min}}^{\textrm{r}}$. The right and left snapshots correspond to the spiral states with counter-clockwise and clockwise turnings, respectively, and the middle one exhibits a non-spiral state. See SM for movies.
In the present paper, we show that transitions between configurational states of the polymer is consistent with a subcritical pitchfork bifurcation. Furthermore, we prove that the diffusion coefficient for the polymer center of mass (CM), $D_{\textrm{CM}}$ is, counterintuitively, a non-monotonic function of the self-propelling (SP) force, $F_{\textrm{sp}}$, acting on different segments of the polymer. Moreover, we find that there exist an optimal self-propelling force $F_{\textrm{sp}}$ that maximizes $D_{\textrm{CM}}$.
Simulation model – We consider a bead-spring semi-flexible polymer with $N=200$ monomers in a two-dimensional, squared simulation box of area $L\!\times\!L$ – where $L=500\sigma$ and $\sigma$ the length unit – with periodic boundary conditions [Fig. 1(a)]. For non-zero value of the bending rigidity coefficient $\kappa_{\textrm{b}}$, the persistence length of the polymer in two dimensions is obtained as $\ell_{\textrm{p}}=2\kappa_{\textrm{b}}/(k_{\textrm{B}}T)$. The $k_{\textrm{B}}T$ is the thermal energy and $k_{\textrm{B}}$ is the Boltzmann constant [see the SM for more details about the simulation model].
To model the active polymer, the polymer is divided into $N/\ell_{\textrm{{p}}}$ segments. Each segment – see Fig. 1(a) – contains a tail (in red color) and a head (in green color) monomers as the first and last monomers, respectively, and the rest of monomers are in blue color. Employing Langevin dynamics (LD) a SP force $F_{\textrm{sp}}$ acts on the head monomer of each segment and in the direction from tail to head in each time step, depicted by the orange arrow and dashed line, respectively in Fig. 1(a) Amir [see the SM for more details]. All the LD simulations are performed using LAMMPS package lammps .
Figure 2: (a) Non-monotonic diffusion coefficient of the polymer center of mass $D_{\textrm{CM}}$ as a function of the SP force $F_{\textrm{sp}}$, from LD simulations (LD Sim; black circles) and the theory obtained from Eq. (4) (Theory; orange squares). (b) The probability of finding the polymer in the non-spiral state $p_{\textrm{ns}}$ (red squares) and in the spiral state $p_{\textrm{s}}$ (blue circles) as a function of the SP force $F_{\textrm{sp}}$. The intersection of the red and blue curves (denoted by the horizontal black dashed line at $p_{\textrm{ns}}=p_{\textrm{s}}=0.5$ and the vertical black dashed-dotted line at $F_{\textrm{sp}}=1.9$) corresponds to the peak of the black curve $D_{\textrm{CM}}$ in panel (a).
Spiral states & transport – As the snapshots in Fig. 1(b) show, the self-propelled polymer possesses different configurations that correspond to spiral and non-spiral states, which are quantified by considering the turning number $\psi$ of the polymer. The turning number of the entire polymer with $N-1$ bonds is defined as  turn_handbook
$\psi=\frac{1}{2\pi}\sum^{N-1}_{i=1}(\theta_{i+1}-\theta_{i}),$ (1)
where $\theta_{i}$ is the angle between the $i^{\textrm{th}}$ bond vector and horizontal axis. Note that $\psi>0$ [right snapshot in Fig. 1(b)] and $\psi<0$ [left snapshot in Fig. 1(b)] imply that the global configuration of the polymer is counter-clockwise (CCW) and clockwise (CW), respectively. A vanishing $\psi$ value indicates either a rod-like configuration or a random coil polymer configuration [middle snapshot in Fig. 1(b)].
In simulations, we find that the probability distribution function (PDF) of the turning number $P(\psi)$ has three peaks: two are located in the left and right hand sides and one at $\psi=0$; see Fig. 1(b) where two vertical black dashed lines, at the two minima of the $P(\psi)$, indicate the region of $\psi$ that corresponds to non-spiral states. The time evolution $\psi(t)$, the PDF of $P(\psi)$ for various values of $F_{\textrm{sp}}$, and different types of transition between the non-spiral and the spiral states have been explained in the SM. In the spiral state, the sum of all active forces acting on different segments of the polymer is almost negligible and thus the polymer barely moves. On the other hand, when the active polymer escapes from a spiral configuration and stays in the non-spiral state, it is less coiled and active forces along the polymer contribute to displace the center of mass. Fig. 2(a) shows the $D_{\textrm{CM}}$ as a function of $F_{\textrm{sp}}$ for a fixed value of $\ell_{\textrm{p}}=5$ (black circles), where it is evident that $D_{\textrm{CM}}$ is non-monotonic with respect to the $F_{\textrm{sp}}$. The typical trajectories of the polymer CM for various values of $F_{\textrm{sp}}$ and the corresponding mean-square-displacement curves have been presented in the SM. Moreover, it exists an optimal value of $F_{\textrm{sp}}$ that maximizes $D_{\textrm{CM}}$. The maximum of $D_{\textrm{CM}}$ takes place when the probabilities of finding the polymer in the spiral and non-spiral states are equal to each other; see Fig. 2(b).
Figure 3: (a) The PDF of the $\psi$, $P(\psi)$ obtained from integration of Eq. (2) with $\chi=-0.184$ corresponding to SP force $F_{\textrm{sp}}=1.9$ and for fixed value of the noise $\Gamma=0.00375$ (orange circles) and from the LD simulations (purple squared). (b) Bifurcation diagram: theory vs. simulation. The blue solid and dashed lines that are obtained from Eq. (3) show the stable and unstable fixed points, respectively. The filled and empty green circles represent the stable and unstable fixed points, respectively, and are coming from the LD simulations. Inset presents the value of $F_{\textrm{sp}}$ as a function of $\chi$. (c) The PDF of $x$, $P(x)$ (top) and the corresponding potential $\Phi(x)$ (bottom). In both panels the value of $\chi$ has been set to $\chi=-0.2$. (d) The probability of finding the polymer in the non-spiral state $p_{\textrm{ns}}$ (red squares from the LD simulation and red solid line from the theory) and in the spiral state $p_{\textrm{s}}$ (blue circles from the LD simulation and blue solid line from the theory) as a function of $\chi$. The intersection of the red and blue curves (denoted by the horizontal black dashed line at $p_{\textrm{ns}}=p_{\textrm{s}}=0.5$ and the vertical black dashed-dotted line at $\chi=-0.184$ corresponds to $F_{\textrm{sp}}=1.9$) corresponds to the peak of the black curve $D_{\textrm{CM}}$ in Fig. 2(a).
These probabilities are defined as $p_{\textrm{s}}=\frac{\langle\tau_{\textrm{s}}\rangle}{\langle\tau_{\textrm{s}}\rangle+\langle\tau_{\textrm{ns}}\rangle}$ and $p_{\textrm{ns}}=\frac{\langle\tau_{\textrm{ns}}\rangle}{\langle\tau_{\textrm{s}}\rangle+\langle\tau_{\textrm{ns}}\rangle}$, where $\langle\tau_{\textrm{s}}\rangle$ and $\langle\tau_{\textrm{ns}}\rangle$ refer to the average time the polymer stays in the spiral and non-spiral states, respectively.
Theory – Next the stability of the turning number is checked by using linear stability analysis strogatz . The PDF of the turning number $\psi$ can be expressed as $P(\psi)\sim\textrm{exp}[-\Phi(\psi)]$ with $\Phi(\psi)=a_{2}\psi^{2}-a_{4}\psi^{4}+a_{6}\psi^{6}$ a potential, where $a_{2},a_{4}$ and $a_{6}$ are constants that depend on the value of $F_{\textrm{sp}}$ and are obtained using the best fitting of the PDF of the turning number coming from LD simulation by the $P(\psi)$ [see Fig. 1(b) and in SM, Fig. 1(b)]; cf. turning . A dynamical system of order $n$th can be quantized by $n$ first order ordinary differential equations strogatz . For the present paper as the turning number $\psi$ is the only dynamical variable of the system, therefore one ordinary differential equation for the $\psi$ in dimensionless units can describe the system as
$\frac{dx}{d\tau}=f(x)+\xi({\tau})=\chi x+x^{3}-x^{5}+\xi({\tau}),$ (2)
where $f(\psi)=-d\Phi(\psi)/d\psi$ is an effective “force”, $x=\psi/U$ and $\tau=t/\mathcal{T}$ with $U^{2}=2a_{4}/(3a_{6})$, $\mathcal{T}=3a_{6}/(8a^{2}_{4})$ and $\chi=-3a_{2}a_{6}/(4a^{2}_{4})$. The values of $a_{2}$, $a_{4}$ and $a_{6}$ as a function of $F_{\textrm{sp}}$ and the corresponding values of $\chi$, $\mathcal{T}$ and $U$ can be found in Table I in the SM. The term $\xi$ refers to a white noise that satisfies $\langle\xi(\tau)\rangle=0$ and $\langle\xi(\tau)\xi(\tau^{{}^{\prime}})\rangle=2\Gamma\delta(\tau-\tau^{{}^{\prime}})$. The noise intensity $\Gamma$ is obtained by mapping the PDF of the turning number in LD simulations with the one resulting from the integration of Eq. (2) [see Fig. 3(a)] and has a fixed value of $\Gamma=0.00375$ for different values of $F_{\textrm{sp}}$. The deterministic form of Eq. (2) – that corresponds to $\Gamma=0$ and thus $\dot{x}=f(x;\chi)$ – undergoes a subcritical pitchfork bifurcation strogatz . Time evolution of the solution of Eq. (2) for different values of the initial normalized turning number $x_{0}$ and for fixed value of $\chi=-0.184$ corresponding to $F_{\textrm{sp}}=1.9$ can be seen in the SM. The bifurcation diagram in the $x-\chi$ plane is shown in Fig. 3(b). The fixed points of Eq. (2) – $f(x^{*};\chi)=0$ – are:
$x^{*}=0,\,\pm\bigg{(}\frac{1\pm\sqrt{1+4\chi}}{2}\bigg{)}^{1/2}\,.$ (3)
For $\chi>0$, $x^{*}=0$ is an (linearly) unstable fixed point since $f^{{}^{\prime}}(x^{*})=\frac{\partial f}{\partial x}(x^{*})\geq 0$. For $\chi\leq 0$, $x^{*}=0$ becomes stable ($f^{{}^{\prime}}(0)<0$) and two unstable fixed points – such that $f^{{}^{\prime}}(x^{*})\geq 0$ – emerge. At $\chi_{\textrm{s}}=-1/4$ there are two additional saddle-node bifurcations [denoted by a black star in the main panel of Fig. 3(b)] which are obtained by considering the condition $1+4\chi_{\textrm{s}}\geq 0$. Note that the fixed points are the extremum values of the $P(\psi)$ [see Fig. 1(b) in SM]. The effect of $x^{5}$ term in Eq. (2) in the bifurcation diagram is to turn around the unstable branches at $\chi=\chi_{\textrm{s}}$ and become stable for $\chi>\chi_{\textrm{s}}$ [solid blue lines in the top and bottom in Fig. 3(b) main panel]. These stable states provide the possibility of jumps and hysteresis as $\chi$ is varied, that is the characteristic of the first-order phase transitions. The filled and empty green circles in Fig. 3(b) correspond to the LD simulations and show the stable and unstable states, respectively. The simulations data and theoretical curves show a perfect agreement. The inset in Fig. 3(b) shows how the values of $F_{\textrm{sp}}$ and $\chi$ relate to each other.
To compute the diffusion coefficient $D_{\textrm{CM}}$, and to put in evidence the run-&-tumble character of the process, we consider that during the spiral states, the sum of all active forces acting on different segments of the polymer is negligible. Thus, in this state, the speed (of the CM) of polymer is $v_{\textrm{s}}\simeq 0$. This implies that we assume that the distance travelled by the polymer in the spiral sates is negligible [see the movie in SM]. On the other hand, in the non-spiral states, the polymer is (in comparison to spiral states) extended, and thus, the sum of active forces is non-zero. In consequence, in the non-spiral state, the polymer moves at speed $v_{\textrm{ns}}>0$, and we assume that it travels an average distance $\ell=v_{\textrm{ns}}\langle\tau_{\textrm{ns}}\rangle$. To determine $v_{\textrm{ns}}$, let us consider that in the non-spiral state, the polymer is an extended rod – that experiences an active force $NF_{\textrm{sp}}/\ell_{\textrm{p}}$ and a friction drag $N\,\gamma$ – and thus, moves at speed $v_{\textrm{ns}}=F_{\textrm{sp}}/\gamma~{}\!\ell_{\textrm{p}}$. All together, this implies that as the polymer goes through a cycle spiral-non-spiral, i.e. involving a transition from the spiral to the non-spiral state, moves an average distance $\ell$. Since the average duration of these cycles is $\Omega=\langle\tau_{\textrm{s}}\rangle+\langle\tau_{\textrm{ns}}\rangle$, in a time $t$, the polymer performs an average of $n=t/\Omega$ of those cycles. It is worth noting that $\langle\tau_{\textrm{ns}}\rangle$ and $\langle\tau_{\textrm{s}}\rangle$ can be estimated from the transition rates to escape from the non-spiral and spiral states considering the associated Kramer’s escape problem that assumes these rates are proportional to the exponential of the potential barrier: $\langle\tau_{\textrm{ns}}\rangle^{-1}\propto\exp(\Delta\Phi_{1}/\Gamma)$ and $\langle\tau_{\textrm{s}}\rangle^{-1}\propto\exp(\Delta\Phi_{2}/\Gamma)$, respectively, where $\Delta\Phi_{1}=\frac{1}{12}(1+4\chi)^{3/2}$ and $\Delta\Phi_{2}=\frac{1}{48}(-1+\sqrt{1+4\chi})(1+8\chi-\sqrt{1+4\chi})$ [see Fig. 3(c) bottom panel]. The detail of how $\Delta\Phi_{1}$ and $\Delta\Phi_{2}$ are obtained, can be seen in the SM. Recall that then $p_{\textrm{s}}$ and $p_{\textrm{ns}}$ can also be expressed in term of these rates, and that $\chi=\chi(F_{\textrm{sp}})$ [see Fig. 3(d)]. Finally, let us recall that for a random walk (RW) after $n_{\textrm{RW}}$ steps of length $\ell_{\textrm{RW}}$, the mean-square displacement goes as $\langle\mathbf{r}^{2}\rangle=n_{\textrm{RW}}\,\ell_{\textrm{RW}}^{2}$. Thus, if we interpret that in time $t$, the $n$ cycles performed by the active polymer correspond to $n_{\textrm{RW}}$ steps of a RW, that moves in each step an average distance $\ell_{\textrm{RW}}=\ell$, we can approximate the diffusion coefficient – using the above-given expressions for $n$ and $\ell$ – as:
$D_{\textrm{CM}}=\lim_{t\to\infty}\frac{\langle\mathbf{r}^{2}\rangle}{4t}=\frac{v_{\textrm{ns}}^{2}}{4}\frac{\langle\tau_{\textrm{ns}}\rangle^{2}}{\langle\tau_{\textrm{s}}\rangle+\langle\tau_{\textrm{ns}}\rangle}=\frac{\overline{v}^{2}\Omega}{4}\,,$ (4)
where $\overline{v}=p_{\textrm{ns}}v_{\textrm{ns}}$, and is in agreement with the run-and-tumble particle model Sokolov_PRE_2012 . The comparison between the diffusion coefficients $D_{\textrm{CM}}$ predicted by Eq. (4) (denoted by ”Theory”; orange squares) and the one obtained from the LD simulations (”LD Sim”; black circles) displayed in Fig. 2(a), shows that there exists a good quantitative agreement.
Importantly, Eqs. (2) and (4) allow us to prove that there is an optimal self-propelling force that maximizes the diffusion coefficient. The existence of an optimal self-propelling force is easy to understand. When $F_{\textrm{sp}}\to 0$, the spiral states is not observed. The speed in the non-spiral state is $v_{0}\propto F_{\textrm{sp}}$, then, $F_{\textrm{sp}}\to 0$ implies $D_{\textrm{CM}}\to 0$. On the contrary, for $F_{\textrm{sp}}\gg 1$, the polymer is locked in spirals, but since we assume that $v_{0}=0$ in this state, we get again  $D_{\textrm{CM}}\to 0$. In summary, increasing $F_{\textrm{sp}}$ increases $v_{0}$, but at the same time enhances the probability of finding the polymer in the spiral state, where the motion of the center of mass of the polymer is strongly reduced. And thus, we find that for intermediate values of $F_{\textrm{sp}}$, $D_{\textrm{CM}}$ is maximized. Compare the results obtained here with those reported in a numerical study with the Krotky-Porod model theo4 .
Our studies puts in evidence the existence of a key coupling between the structural dynamics of the polymer and its transport properties. Moreover, we have shown that there exist an optimal active force $F_{\textrm{sp}}$ that maximizes the diffusion coefficient $D_{\textrm{CM}}$. The reported results are of relevance for a large number of experimental active polymer studies, including actin filament motion in motility assays experiments and DNA duplication, among many other examples.
Acknowledgements.
FP acknowledges fruitful discussions with Debasish Chaudhury.